OpenStudy (3psilon):

Three blocks of masses 1.0, 2.0, and 4.0 kilograms are connected by massless strings, one of which passes over a frictionless pulley of negligible mass, as shown below. Calculate The acceleration of the 4.0 kilogram block.

OpenStudy (3psilon):

There's also an image

OpenStudy (anonymous):

4g-T=4a T-3g=3a 4g-(3g+3a) = 4a g=7a a=g/7

OpenStudy (3psilon):

@Algebraic! Where did you get T - 3g = 3a

OpenStudy (3psilon):

The acceleration and tension is the same on the left?

OpenStudy (anonymous):



OpenStudy (anonymous):

sum of forces = ma

OpenStudy (anonymous):

T-3g =3a

OpenStudy (anonymous):

try to do all your pulley problems like that. much easier when direction is always consistent.

OpenStudy (anonymous):